OpenStudy (madeleine2):

Write the standard form of the line that passes through the given points. (6, 1) and (5, 4)

OpenStudy (mathmale):

Please use the slope formula to obtain the slope of the straight line that passes through (6,1) and (5,4). "Standard form of the equation of a line" looks like Ax + By + C = 0. What else do you need to know to complete the solution of this problem?

563blackghost (563blackghost):

First determine the slope. \(\huge\bf{\frac{y_{2}-y_{1}}{x_{2}-x_{1}} \rightarrow \frac{4-1}{5-6}}\)

Directrix (directrix):

@madeleine2 4 - 1 = ?